### v1

# The Primacy of Magnitude in Low-Rank Adaptation

Abstract
Low-Rank Adaptation (LoRA) offers a parameter-efficient paradigm for tuning large models. While recent spectral initialization methods improve convergence and performance over the naive “Noise & Zeros” scheme, their extra computational and storage overhead undermines efficiency. In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM [^1] Code will be released at this page, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies. Our key contributions are threefold: (i) Magnitude of weight updates determines convergence. We prove low-rank structures intrinsically bound update magnitudes, unifying hyperparameter tuning in learning rate, scaling factor, and initialization as mechanisms to optimize magnitude regulation. (ii) Spectral initialization succeeds via magnitude amplification. We demystify that the presumed knowledge-driven benefit of the spectral component essentially arises from the boost in the weight update magnitude. (iii) A novel and compact initialization strategy, LoRAM, scales deterministic orthogonal bases using pretrained weight magnitudes to simulate spectral gains. Extensive experiments show that LoRAM serves as a strong baseline, retaining the full efficiency of LoRA while matching or outperforming spectral initialization across benchmarks.
1 Introduction
The rise of large pretrained models Brown2020LanguageMA ; OpenAI2023GPT4TR ; touvron2023llama ; bai2023qwen ; guo2025deepseek has driven urgent needs for parameter-efficient fine-tuning (PEFT) methods peft ; Hu2021LoRALA ; Lester2021ThePO ; He2021TowardsAU ; Edalati2022KronAPE ; zhang2025parameter . Among these, Low-Rank Adaptation (LoRA) Hu2021LoRALA stands out for its efficiency, flexibility, and stability. By freezing pretrained weights and injecting trainable low-rank matrices, LoRA enables the update of less than 1% of the parameters, significantly reducing memory and compute costs. Its plug-and-play nature achieves easy integration into diverse models, facilitating model sharing and federated learning SLoRA ; FeDeRA . Furthermore, LoRA helps prevent catastrophic forgetting biderman2024lora , making it suitable for continuous learning. These advantages have led to its wide adoption in NLP liu2022few ; Ding2022DeltaTA ; sun2023comparative ; zhao2024lora and multimodal applications ye2023mplug ; guo2023animatediff ; blattmann2023stable ; ruiz2023dreambooth ; zhang2023towards .
Despite achieved efficiency, the low-rank reparameterization constrains practical performance and convergence biderman2024lora . Besides the well-known “representation bottleneck” hyeon2021fedpara ; Edalati2022KronAPE ; Zi2023DeltaLoRAFH ; gao2024parameter ; jiang2024mora ; lialin2023relora ; xia2024chain ; PeriodicLoRA , LoRA is highly sensitive to hyperparameters due to the non-convex and non-smooth loss landscape li2024crucial . Effective training relies on careful tuning of rank zeng2023expressive ; Zhang2023AdaptiveBA ; ding2023sparse , scaling factor kalajdzievski2023rank , learning rate lora+ , initialization strategies zhu2024asymmetry ; hayou2024impact ; li2024crucial , and preconditioning li2024crucial ; zhang2024riemannian ; lora-one . Recent works increasingly leverage information from pretrained weights PiSSA ; MiLoRA ; olora or task-specific data corda ; lora-ga ; lora-one to improve the “Noise & Zeros” baseline. Among these, PiSSA PiSSA pioneers the use of Singular Value Decomposition (SVD) for LoRA initialization, employing spectral components of pretrained weights to significantly enhance convergence and performance. Subsequent works MiLoRA ; lora-ga ; olora ; lora-one ; corda extend to diverse matrix decompositions, fostering a wave of knowledge-driven initialization schemes.
While spectral initialization PiSSA ; MiLoRA ; lora-ga ; lora-one ; corda showcases considerable promise in convergence and performance, two fundamental challenges persist. First, they introduce complexity by requiring additional matrix decomposition and storage overhead, undermining usage efficiency in resource-constrained settings vera ; gao2024parameter ; QLoRA and hindering seamless integration with deep learning libraries peft ; wolf-etal-2020-transformers . Second, their success remains poorly understood. The common justification PiSSA ; MiLoRA ; corda ; mao2025survey that spectral components preserve features better than random alternatives lacks theoretical grounding. Although recent works lora-ga ; lora-one suggest that some specific initialization may approximate full-parameter gradients, the non-convex nature of LoRA renders training dynamics unpredictable.
In this paper, we demystify the knowledge-driven intuition behind spectral initialization and reveal that its effectiveness primarily stems from the magnitude scaling of weight updates. We design a minimal “Basis & Basis” initialization strategy, which demonstrates comparable performance without the overhead of SVD operations. Specifically, our key contributions can be summarized as:
  • We identify weight update magnitude as a fundamental principle for analyzing and improving LoRA’s training dynamics. This principle unifies previously independent factors, such as learning rate, scaling, and initialization, revealing their shared ability to control weight update strength and achieve comparable amplification effects when properly configured.
  • We demonstrate that initialization scheme critically shapes LoRA’s weight magnitude dynamics. Theoretically, we prove that LoRA naturally produces smaller updates than full fine-tuning, which limits its convergence and expressiveness. Moreover, we show that spectral initialization amplifies updates, providing a principled explanation for its effectiveness beyond knowledge-driven intuition.
  • Guided by the magnitude principle, we propose Magnitude-driven Initialization (LoRAM) to make LoRA initialization efficient again. LoRAM employs a logarithmic magnitude factor to retain the benefits of spectral scaling, while directly scaling deterministic orthogonal bases to eliminate the need for decomposition and storage. Extensive experiments on language and vision-language tasks establish LoRAM as a strong and practical baseline, surpassing prior initialization schemes.
Figure 1: We propose LoRAM, a magnitude-driven initialization method that enhances both the convergence and performance of LoRA while maintaining its efficiency. Unlike spectral initialization, which precomputes and stores singular components ($U,V,S$) PiSSA , LoRAM uses deterministic orthogonal bases and derives scaling from pretrained weight statistics. This elegant simplification is grounded in our analysis of LoRA through a novel lens of magnitude dynamics, where we show that the benefits of spectral values in scaling weight update magnitude can be effectively approximated.
2 Magnitude Principle for Characterizing LoRA Dynamics
2.1 Preliminaries and Notations
Given a pretrained weight matrix $W\in\mathbb{R}^{n\times m}$, LoRA Hu2021LoRALA reparameterizes the forward pass as
where $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ are trainable low-rank matrices with $r\ll\min(n,m)$, and $\alpha$ scales the update magnitude. The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for computation convenience, as illustrated in Figure 1(b). Given a loss function $L$, LoRA updates are computed as:
$\nabla_{A}L=\alpha B^{\top}\frac{\partial L}{\partial y}x^{\top}=\alpha B^{\top}(\nabla_{W}L),\quad\nabla_{B}L=\alpha\frac{\partial L}{\partial y}x^{\top}A^{\top}=\alpha(\nabla_{W}L)A^{\top}.$
Magnitude metric. To elucidate how LoRA affects the training process, we analyze the dynamics of the parameters $A$, $B$, and the resulting weight update $W_{\text{LoRA}}$. Specifically, we define the weight magnitude as $\nu[W_{\text{LoRA}}]=\frac{1}{mn}\|W_{\text{LoRA}}\|_{F}^{2}$, which serves as a central metric in our study. Assuming independent and zero-mean entries he2015delving , the expected weight magnitude is given by $\mathbb{E}[\nu[BA]]=r\,\mathbb{E}[\nu[B]]\,\mathbb{E}[\nu[A]]$. In the asymptotic regime where $m$ and $n$ are large, $\nu[W_{\text{LoRA}}]$ is approximated with the variance of $W_{\text{LoRA}}$, and $\nu[BA]\approx r\,\nu[B]\,\nu[A]$. Our analysis is motivated by the fact that LoRA introduces no change to the pretrained weights initially, i.e., $\nu[W^{(0)}_{\text{LoRA}}]=0$, while its effect emerges gradually through training. Therefore, we use the term “magnitude” instead of “variance” to highlight the cumulative growth of $W_{\text{LoRA}}$. In Appendix B, we also present a theoretical insight showing that parameter magnitude is a key determinant of LoRA’s expressiveness.
2.2 Effect of Hyperparameters on Update Magnitude
Let $\Delta W^{(t)}_{\mathrm{LoRA}}$ denote the weight update at step $t$ for LoRA framework, and $W^{(t)}_{\text{LoRA}}=\sum_{i=0}^{t-1}\Delta W^{(i)}_{\mathrm{LoRA}}$ represent cumulative adaptation. Given a learning rate $\eta$, we expand $\Delta W^{(t)}_{\mathrm{LoRA}}$ using gradient update rules, leading to the following formulations for the update magnitude [^1] See Appendix for derivation and proof:
$\Delta W^{(t)}_{\mathrm{LoRA}}$ $=\alpha\eta\left(B^{(t)}\nabla_{A}L^{(t)}+\nabla_{B}L^{(t)}A^{(t)}+\eta\nabla_{B}L^{(t)}\nabla_{A}L^{(t)}\right),$ (3)
$\text{and}\quad\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]$ $\approx r\alpha^{2}\eta^{2}\left(\nu[B^{(t)}]\nu[\nabla_{A}L^{(t)}]+\nu[\nabla_{B}L^{(t)}]\nu[A^{(t)}]\right).$ (4)
These equations indicate the complex interplay of multiple hyperparameters, distinct from the full-parameter updates given by $\Delta W^{(t)}=-\eta\nabla_{W}L^{(t)}$. We investigate the interplay among the learning rate $\eta$, scaling factor $\alpha$, and initialization magnitude, revealing a quantifiable equivalence relationship.
Proposition 1 (Parameter Scaling Equivalence).
For LoRA layers defined in Eq. (1), consider decomposing the scaling factor $\alpha=\alpha^{\prime}\alpha_{A}\alpha_{B}$, where $\alpha^{\prime},\alpha_{A},\alpha_{B}\in\mathbb{R}^{+}$. Under the commonly used optimization frameworks with negligible numerical errors, the following parametrization schemes exhibit dynamical equivalence throughout training: For all iterations $t\geq 0$, $\Delta W^{(t)}_{\mathrm{LoRA}}=\Delta\tilde{W}^{(t)}_{\mathrm{LoRA}}$ and $W^{(t)}_{\mathrm{LoRA}}=\tilde{W}^{(t)}_{\mathrm{LoRA}}$, where $\tilde{A}^{(t)}$, $\tilde{B}^{(t)}$ and $\tilde{W}^{(t)}$ represent the re-parameterized versions.
Original SGD Adam
Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
Remarks. This equivalence underscores how hyperparameters collectively regulate update magnitude, effectively reducing the search space for optimal configurations. A striking implication is that, increasing $\eta_{B}$ in LoRA+ lora+ is identical to scaling $\alpha$ in RsLoRA kalajdzievski2023rank under the "Noise & Zeros" initialization, highlighting the critical role of initialization magnitude in shaping LoRA’s training dynamics, which is rarely discussed in prior works. For the non-zero initialization, it is advisable to first adjust the initialization magnitudes and learning rates for moderate improvements, as modifying $\alpha$ inherently combines the effects of both and may result in drastic and unpredictable changes.
To demonstrate the joint effect of hyperparameters on LoRA dynamics, we conduct a controlled experiment using a 5-layer MLP with “Noise & Zeros” initialized LoRA layers, setting the intermediate dimension to 400 and the LoRA rank to 25. The network is trained on synthetic data under various hyperparameter settings, consequently using SGD and Adam optimizers with $\eta=5\times 10^{-5}$. As shown in Figure 2(a), all settings with $\alpha=16$ result in identical loss trajectories and parameter evolution, confirming the theoretical predictions. In contrast, weight updates deviate significantly when using $\alpha=1$ with $\eta$ scaled by 4, indicating that equivalence holds only under specific rules.
Figure 2: (a) Validation of Proposition 1. Each curve represents a model with unique hyperparameters. The norm difference (right axis) aggregates Frobenius norm discrepancies between the baseline model (black) and others across layers. Purple and other curves share identical learning rates but diverge due to differing initialization magnitudes. Equivalent optimization trajectories emerge from diverse hyperparameter combinations under both SGD and Adam optimizers. (b) Validation of Proposition 2. The black curve represents random orthogonal initialization. Parameter magnitudes are predominantly governed by initialization scaling, resulting in smaller weight changes compared to conventional linear layers. This necessitates the magnitude scaling in enhancing LoRA performance.
2.3 Magnitude Limitation Rooted in Low-Rank Structure
Guided by the established equivalence framework, we fix $\alpha=1$ to eliminate the interference of scaling factors, which is the most commonly-used configuration in practical implementation. In the following, we prove initialization magnitudes and other factors critically influence the weight update.
Proposition 2 (Parameter Magnitude Dynamics).
Consider LoRA parameters updated with the same learning rate $\eta$. Assume: $A^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{A}^{2}I)$, $B^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{B}^{2}I)$, $\nabla_{W}L^{(t)}\sim\mathcal{N}(\mathbf{0},\sigma_{L}^{2}I)$, and $\mathbb{E}[\langle A^{(t)},\nabla_{A}L^{(t)}\rangle]=\mathbb{E}[\langle B^{(t)},\nabla_{B}L^{(t)}\rangle]=0$. Under these conditions, the parameter magnitudes $\boldsymbol{\nu}_{t}=\left[\mathbb{E}[\nu[A^{(t)}]],\mathbb{E}[\nu[B^{(t)}]]\right]^{T}$ evolve as a linear dynamical system. Its exponential solution admits the linearized approximation under small-$\eta$ regime:
$\boldsymbol{\nu}_{t}=\left(I+\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{t-1}\approx\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix},$
where $\gamma_{A}=m\eta^{2}\sigma_{L}^{2}$, $\gamma_{B}=n\eta^{2}\sigma_{L}^{2}$. This further yields the evolution of weight update magnitude:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
Remarks.
This analysis uncovers essential properties of LoRA initialization and optimization trajectory. First, since $\gamma_{A}$ and $\gamma_{B}$ are very small values ($\ll 1$), the magnitudes of parameters $A^{(t)}$ and $B^{(t)}$ remains nearly unchanged throughout training, potentially constraining the representation capacity of the learned model. Moreover, unlike full-parameter tuning, which evolves at a linear rate of $\gamma$, the low-rank structure introduces a proportional factor $k_{1}$, significantly slowing updates. For instance, the naive “Noise & Zeros” initialization yields $k_{1}=\frac{r}{m}$, while the dimension $m$ in large models like LLaMA touvron2023llama is in the thousands or more. Despite the quadratic term accelerates growth, small gradients may temper this effect in the later training stages.
Figure 2(b) visualizes LoRA magnitude dynamics during training. We use the same network as in Figure 2(a) but apply a nonzero initialization with $\sigma_{A}=\sigma_{B}=\frac{1}{20}$. We explore a regular MLP with the same magnitude denoted as “Linear”, and a group of networks with larger initialization magnitudes. Notably, while the loss decreases significantly, the magnitudes of $A$ and $B$ remain nearly unchanged throughout training. The magnitude evolution of $W$ reveals that the basic LoRA with theoretically $k_{1}=\frac{1}{16}$, grows substantially slower than the regular MLP. Increasing the initialization magnitude effectively accelerates the growth of $W$, aligning with our theoretical analysis.
Integrating analyses in this section, we derive a magnitude principle for LoRA development:
A valid improvement to LoRA convergence will enhance weight update magnitude $\nu[W_{\text{LoRA}}]$.
As shown in Proposition 2, magnitude principle could unify and explain improvement factors in existing works, including the learning rate, scaling factor, gradients, rank and initialization schemes.
3 Demystifying Spectral Gains with Magnitude Principle
The sheer scale of modern neural networks complicates the determination of optimal LoRA initialization across layers. Drawing inspiration from recent spectral initialization methods PiSSA ; MiLoRA ; corda ; lora-ga ; lora-one , which have demonstrated improved convergence and task performance, we reinterpret their effectiveness through the lens of the magnitude principle and introduce a magnitude-driven initialization method called LoRAM. Notably, these methods requires extra SVD computations and storage, leading to increased resource overhead and implementation complexity. In contrast, LoRAM mitigates these drawbacks and achieves even better performance. In the following, we take the seminal work PiSSA PiSSA as a representative baseline and ablate other methods in experiments (see Section 4.3).
3.1 Magnitude Gain in Spectral Initialization
The PiSSA method PiSSA initializes LoRA using the spectral decomposition of pretrained weight matrix $W=USV^{\top}$, which has a rank of $\mathcal{R}[W]$. The spectral initialization is defined as:
$A^{(0)}=A_{\mathrm{SVD}}=\sqrt{S_{r}}V_{:,:r}^{\top},\quad B^{(0)}=B_{\mathrm{SVD}}=U_{:,:r}\sqrt{S_{r}},$
where $S_{r}\in\mathbb{R}^{r\times r}$ contains the top-$r$ singular values, and $U\in\mathbb{R}^{n\times n}$, $V\in\mathbb{R}^{m\times m}$ are the left and right singular vector matrices. While prior works PiSSA ; MiLoRA ; corda ; mao2025survey attribute PiSSA’s success to its ability to preserve principal components, we show that its key advantage lies in singular value weighting. By redistributing dominant variance into the initialization, PiSSA facilitates adaptive magnitude updates across layers, accelerating convergence.
Consider the statistics of the top-$r$ singular values, we define the spectral concentration factor as:
$\rho[r]\triangleq\frac{\mathbb{E}_{r}[s]^{2}}{\mathbb{E}_{\mathcal{R}[W]}[s^{2}]}=\frac{\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}}{\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}}.$
This factor captures the concentration of energy in the top-$r$ singular values. We then reformulate
$\nu(A_{\mathrm{SVD}})=\mathbb{E}_{r}[s]\nu[V_{:,:r}]=\sqrt{\frac{n\rho[r]\nu[W]}{m\mathcal{R}[W]}},\ \nu(B_{\mathrm{SVD}})=\mathbb{E}_{r}[s]\nu[U_{:,:r}]=\sqrt{\frac{m\rho[r]\nu[W]}{n\mathcal{R}[W]}}.$
Essentially, $\rho[r]$ acts as a scaling factor that redistributes variance from the pretrained weight matrix, influencing the magnitude of updates during training. Since $\rho[r]$ monotonically decreases with $r$, its impact is more pronounced for smaller $r$, making it particularly relevant for LoRA applications.
Figure 3: Illustration of spectral gain factor $Q[r]$ defined in Eq. (10) and spectral concentration factor $\rho[r]$ defined in Eq. (8) across DeBERTa-v3-base he2021debertav3 , LLaMA-2-7B touvron2023llama and FLUX.1-12B flux2024 . Values are computed from uniformly sampled layers. The white dotted line represents the linear growth rate of naive LoRA weight magnitudes, while spectral initialization exhibits faster growth. Due to its concave nature, we approximate the spectral gain factor using a logarithmic function.
Spectral Gain Factor. Taking the above magnitudes into the dynamics in Eq. (6) further derives:
$k_{1}=Q[r](m+n)\nu[W],\quad 0\leq Q[r]\triangleq\frac{\rho[r]r}{\mathcal{R}[W]}\leq 1.$
Given that $\nu[W]\sim\mathcal{O}(\min(\frac{1}{m},\frac{1}{n}))$, this results in a gain factor of at least $Q[r]$, which we term as the “spectral gain factor”. As shown in Figure 3, the spectral gain factor $Q[r]$ exhibits bounded variation in $[0,1]$ with characteristic concavity, which can be formally derived via Jensen’s inequality. Although SVD components are not completely independent preventing the theoretical monotonic increase in $Q[r]$, this concavity also suggests that the gain effect is more pronounced when $r$ is small, reinforcing the effectiveness of spectral initialization for LoRA in the parameter-efficient manner.
3.2 Efficient Magnitude-driven Initialization with LoRAM
We propose LoRAM to achieve similar magnitude update rate in Eq. (10) like PiSSA while eliminating spectral computation. As depicted in Algorithm 1, LoRAM initializes the parameter matrices as:
$A^{(0)}=A_{\mathrm{LoRAM}}=\beta\cdot\Phi_{m}^{\top},\quad B^{(0)}=B_{\mathrm{LoRAM}}=\beta\cdot\Phi_{n},\quad\beta=\left(\frac{Q[r]\cdot\nu[W]}{\nu\left[\Phi_{n}\Phi_{m}^{\top}\right]}\right)^{\frac{1}{4}}.$
Here $\Phi_{n}$ and $\Phi_{m}$ denote the first $r$ columns of an $n$- and $m$-dimensional orthogonal basis matrices, respectively. Given that $\nu[\Phi_{n}]=\frac{1}{n}$ and $\nu[\Phi_{m}]=\frac{1}{m}$, LoRAM achieves the similar magnitude as PiSSA with $k_{1}=Q[r](m+n)\nu[W].$ We can also derive $\nu[B^{(0)}A^{(0)}]=Q[r]\nu[W]$, implying LoRAM inherently ensures numerical stability and moderate corrections to the pretrained weight.
Logarithmic Gain Factor. Due to the concave nature of $Q[r]$, we approximate its analytical form using an asymptotic expansion: $Q[r]\approx\log_{\min(n,m)}(r).$ As illustrated in Figure 3, this logarithmic function effectively captures the monotonic increase nature, providing a predictable improvement than LoRA and PiSSA particularly when using a small rank $r$.
Deterministic Basis Matrix. To eliminate the need for storing initialization buffers, we adopt an analytic approach instead of random initialization. Specifically, we employ the Discrete Sine Transform (DST) basis due to its simplistic mathematical definition:
$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 1\leq i,j\leq m.$
This formulation constructs orthogonal matrices of arbitrary dimensions, ensuring reproducibility across different devices while providing provable statistical properties: $\mathbb{E}[\Phi_{m}]=0$ and $\nu[\Phi_{m}]=\frac{1}{m}$. One may wonder if randomness is required for initialization, we find it unnecessary in LoRA. In fact, DST even slightly outperforms random strategies in our ablation experiments (see Section 4.3).
Efficiency and Compatibility. Since $\beta$ and $\Phi$ avoid complex matrix operations, LoRAM retains the efficiency and storage footprint of naive LoRA. As it only modifies initialization, LoRAM remains plug-and-play, integrating seamlessly into any pipeline that supports standard LoRA. This is especially valuable for modern large models built on fixed and highly optimized frameworks. In contrast, other initialization methods require costly preprocessing, such as matrix generation lora-ga ; corda or decomposition PiSSA ; olora ; MiLoRA , which complicates adoption in standard workflows.
Input: Pretrained weight $W\in\mathbb{R}^{n\times m}$, target rank $r$
Output: Initialized parameters $A^{(0)},B^{(0)},W$
$\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r)
  // Generate basis matrices, Eq (12)
$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$
  // Compute magnitude gain factor
$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$
  // Initialize parameters
Algorithm 1 LoRAM Initialization Procedure
4 Experiments
We conduct comprehensive experiments to evaluate LoRAM efficiently implemented via the PEFT library peft . Following conventional settings PiSSA ; MiLoRA ; lora+ , we assess performance on language tasks and extend the evaluation to vision-language tasks, demonstrating LoRAM’s generalization across diverse models and modalities. All experiments are run on servers with 8 NVIDIA H800 GPUs.
Baselines. While extensive research on LoRA has explored aspects like structural modifications and rank control, these directions are largely orthogonal to our focus on hyperparameter analysis within the naive LoRA framework. In line with this, we compare LoRAM with the naive LoRA (ICLR 2022) Hu2021LoRALA , as well as several representative hyperparameter tuning strategies. We first consider weight-driven initialization (marked “§$\S$”), including PiSSA (NeurIPS 2024) PiSSA , which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA , which utilizes the last $r$ singular vectors and values; and OLoRA olora , which applies orthogonal initialization via QR decomposition. All these methods adopt a fixed scaling factor $\alpha=1$. We then include RsLoRA kalajdzievski2023rank  (marked $\dagger$”), which enhances performance by setting $\alpha=\sqrt{r}$, and LoRA+ (ICML 2024) lora+  (marked $\ddagger$”), which increases the learning rate with the recommended $\eta_{B}=4\eta_{A}$. We also evaluate data-driven initialization in the ablation study, including LoRA-GA (NeurIPS 2024) wang2024lora and CorDA (NeurIPS 2024) corda , which require extra pipeline to leverage training data information. Most of these baselines have been integrated and validated in the PEFT library.
Table 1: Comparison of LoRAM versus hyperparameter tuning baselines on NLG tasks. Experiments conducted with LLaMA2-7B model using two ranks, reporting mean $\pm$ std results (%) over three runs. Bold and underlined values represent the best and second-best performances, respectively.
Rank ##\##Param Method GSM8K MATH HumanEval MBPP Commonsense
N/A 6738M Full FT 60.34 $\pm$ 1.32 11.74 $\pm$ 0.63 32.30 $\pm$ 1.26 39.27 $\pm$ 1.01 79.20 $\pm$ 1.20
16 40M LoRA 31.51 $\pm$ 0.31 4.16 $\pm$ 0.27 15.98 $\pm$ 0.20 28.65 $\pm$ 0.47 66.56 $\pm$ 1.21
RsLoRA 39.04 $\pm$ 0.53 4.94 $\pm$ 0.40 18.85 $\pm$ 0.66 28.10 $\pm$ 0.64 73.24 $\pm$ 0.84
LoRA+ 31.69 $\pm$ 0.64 3.98 $\pm$ 0.38 18.54 $\pm$ 0.52 28.00 $\pm$ 0.81 72.19 $\pm$ 1.43
MiLoRA§ 29.70 $\pm$ 0.42 4.18 $\pm$ 0.21 14.69 $\pm$ 0.66 27.23 $\pm$ 0.53 67.90 $\pm$ 1.20
OLoRA§ 35.83 $\pm$ 0.58 4.80 $\pm$ 0.53 16.58 $\pm$ 0.38 27.44 $\pm$ 0.76 73.48 $\pm$ 1.09
PiSSA§ 37.68 $\pm$ 0.45 5.16 $\pm$ 0.41 18.37 $\pm$ 0.49 28.62 $\pm$ 0.68 73.72 $\pm$ 1.05
LoRAM§ 40.32 $\pm$ 0.43 5.30 $\pm$ 0.37 18.92 $\pm$ 0.55 28.83 $\pm$ 0.63 75.19 $\pm$ 1.10
128 320M LoRA 40.27 $\pm$ 0.70 4.72 $\pm$ 0.43 20.11 $\pm$ 0.32 28.84 $\pm$ 0.37 73.64 $\pm$ 1.13
RsLoRA 50.38 $\pm$ 0.37 7.32 $\pm$ 0.28 21.32 $\pm$ 0.70 30.73 $\pm$ 0.43 77.01 $\pm$ 1.17
LoRA+ 40.41 $\pm$ 0.67 5.28 $\pm$ 0.53 20.71 $\pm$ 0.88 29.13 $\pm$ 0.78 78.19 $\pm$ 1.33
MiLoRA§ 39.81 $\pm$ 0.89 5.18 $\pm$ 0.58 20.39 $\pm$ 0.21 29.95 $\pm$ 1.05 74.29 $\pm$ 1.09
OLoRA§ 50.10 $\pm$ 0.64 7.01 $\pm$ 0.56 20.72 $\pm$ 0.67 30.21 $\pm$ 0.89 78.61 $\pm$ 0.97
PiSSA§ 51.48 $\pm$ 0.77 7.04 $\pm$ 0.54 21.62 $\pm$ 0.48 31.07 $\pm$ 0.68 77.28 $\pm$ 0.98
LoRAM§ 51.12 $\pm$ 0.73 7.25 $\pm$ 0.68 22.03 $\pm$ 0.56 31.53 $\pm$ 0.72 77.81 $\pm$ 0.96
Table 2: Comparison of LoRAM versus hyperparameter tuning baselines on GLUE benchmark. Experiments conducted with the DeBERTa-v3-base model using rank 8, reporting mean results over three runs. Bold and underlined values represent the best and second-best performances, respectively.
Method ##\##Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Full FT 184M 88.31 93.57 89.46 67.26 92.80 91.52 83.75 86.87
LoRA 1.33M 90.23 95.87 84.06 63.56 93.88 90.55 50.18 87.20
RsLoRA 1.33M 90.33 95.64 86.38 64.85 93.97 90.26 60.31 88.37
LoRA+ 1.33M 90.37 95.32 87.54 64.79 94.32 90.93 65.37 89.20
MiLoRA§ 1.33M 90.28 95.75 87.00 62.58 93.97 90.83 54.87 87.74
OLoRA§ 1.33M 90.19 94.83 88.72 65.59 93.36 90.74 74.09 88.52
PiSSA§ 1.33M 90.38 95.64 89.21 65.06 93.84 91.35 74.36 88.90
LoRAM§ 1.33M 90.3490.34{90.34}90.34 95.29 89.95 65.53 94.08 91.70 74.72 89.93
Table 3: Comparing LoRAM with other LoRA variants on LLaVA for multimodel tasks. Bold and underlined values indicate the top and second-best performances, respectively.
Method $\textbf{MME}_{\text{Cog}}$ $\textbf{MME}_{\text{Per}}$ MMMU AI2D ChartQA OCRBench TextVQA ScienceQA
Full FT 280 1541 0.355 0.583 0.251 0.361 0.597 0.722
LoRA 278 1402 0.331 0.557 0.231 0.333 0.536 0.684
RsLoRA 274 1385 0.334 0.573 0.227 0.328 0.539 0.694
LoRA+ 283 1389 0.341 0.565 0.229 0.331 0.545 0.690
MiLoRA§$\S$ 285 1354 0.340 0.564 0.220 0.335 0.536 0.681
OLoRA§$\S$ 288 1404 0.345 0.565 0.228 0.330 0.540 0.677
PiSSA§$\S$ 311 1411 0.344 0.564 0.232 0.338 0.547 0.686
LoRAM§$\S$ $\underline{308}$ 1406 0.350 0.571 0.238 0.336 0.551 0.700
Figure 4: Comparison of LoRA, PiSSA, and LoRAM on image customization task. Experiments conducted with the state-of-the-art FLUX.1-12B model using rank 8.
4.1 Evaluating the Performance on Natural Language Tasks
Nature Language Generation (NLG). As shown in Table 1, we conduct supervised fine-tuning of LLaMA 2-7B touvron2023llama on math, coding, and commonsense reasoning tasks. Our setup strictly follows PiSSA PiSSA , using the AdamW optimizer with a batch size of 128, a learning rate of $2\times 10^{-5}$, a warmup ratio of 0.03, and no weight decay. All experiments are performed on subsets containing 100K data points for one epoch to minimize training overhead. For math tasks, the model is tuned on MetaMathQA yu2023metamath and evaluated on GSM8K cobbe2021gsm8k and MATH yu2023metamath validation sets. For coding tasks, we use CodeFeedback zheng2024opencodeinterpreter as training dataset, with evaluations on HumanEval chen2021evaluating and MBPP austin2021program . For commonsense tasks, model is tuned on Commonsense170K hu2023llmadapters , and we report averaged accuracy on eight sub-datasets. The results in Table 1 demonstrate that LoRAM consistently outperforms LoRA variants across diverse tasks and rank settings, without requiring matrix decomposition.
Nature Language Understanding (NLU). We evaluate the NLU performance by fine-tuning the DeBERTa-v3-base model he2021debertav3 with a rank of 8 on eight tasks in the GLUE benchmark Wang2018GLUEAM . We utilize scripts from the Transformers Library wolf-etal-2020-transformers to ensure a fair comparison. All methods are trained with a learning rate of $1\times 10^{-4}$ for 3 training epochs, except for MRPC, which uses 5 epochs due to its smaller size. We report overall matched and mismatched accuracy on MNLI, Matthew’s correlation on CoLA, Pearson correlation on STS-B, and accuracy on the other datasets. As shown in Table 2, LoRAM achieves competitive performance against PiSSA across most tasks.
4.2 Evaluating the Performance on Vision-Language Tasks
Text-to-image synthesis. We adapt the advanced FLUX.1-12B flux2024 to address the image customization task, implementing LoRA, PiSSA, and LoRAM under identical configurations: a learning rate of $1\times 10^{-4}$, a batch size of 1 and 1,000 iterations. We set the rank to 8, optimizing 9.3 million parameters while maintaining computational efficiency on a single GPU. The training data and prompt template adhere to DreamBooth’s protocol ruiz2023dreambooth . Qualitative results in Figure 4 demonstrate that LoRAM exhibits marginally superior performance in detail fidelity compared to PiSSA.
Image-to-text generation. Following the pipeline of LLaVA LLaVA , we employ CLIP-ViT-L/14 radford2021learning as the visual encoder, Vicuna-13B zheng2023judging as the text decoder, and a new visual resampler guo2024llava as the connector. In the pre-training stage, we fine-tune only the perceiver resampler using the CC-595K dataset LLaVA for one epoch. During the subsequent instruction-tuning stage, we fine-tune both Vicuna and the resampler using a 656K mixture dataset LLaVA . The learning rate is set to $2\times 10^{-5}$, and the batch size is 128. We follow the official implementation and use a rank of $64$. As shown in Table 3, LoRAM achieves favorable performance across multiple multimodal benchmarks.
Training curves. We provide representative training loss curves of diverse initialization methods in Figure 5. It can be noticed that LoRAM is able to have a faster convergence rate in the early stages compared to other LoRA variants and incur smaller losses in the end.
Figure 5: Illustration of training loss curves. LoRAM achieves comparative convergence dynamics to PiSSA across diverse models and benchmarks. See tables and texts for the evaluation results.
Table 4: Results of ablation study. “Orth.” denotes the random orthogonal initialization. The left and right sides of the slash indicate the results of the method and the tracking mode, respectively. The average value is calculated over all the ranks and tasks to compare the overall trend of change.
Rank Task $\mathbf{Q[r]}$ Basis Weight-driven Data-driven + RsLoRA $\log\frac{r}{2}$ $\log r^{*}$ $\log 2r$ Orth. Gaussian MiLoRA PiSSA CorDA LoRA-GA LoRA-GA LoRAM 16 MATH 5.08 5.30 5.18 4.74 4.62 4.18 / 4.05 5.16 / 5.10 4.60 / 4.14 5.73 / 3.76 7.94 7.22 GSM8k 40.1 40.3 40.7 36.3 35.8 29.7 / 29.6 37.6 / 36.7 36.2 / 30.7 45.7 / 30.9 51.5 52.1 MBPP 28.8 28.8 28.3 28.6 27.5 27.2 / 27.8 28.6 / 29.1 25.7 / 28.3 28.3 / 27.8 33.9 31.5 HumanEval 17.1 18.9 17.1 17.7 17.7 14.6 / 14.7 18.3 / 17.3 15.2 / 15.2 19.5 / 17.7 22.0 18.3 128 MATH 7.52 7.25 7.51 7.40 7.62 5.18 / 5.04 7.04 / 6.95 6.26 / 5.04 9.18 / 7.32 9.08 11.1 GSM8k 50.7 51.1 50.4 50.2 49.8 39.8 / 39.2 51.4 / 49.5 44.5 / 40.3 54.4 / 50.8 53.6 59.4 MBPP 31.5 31.5 32.3 31.3 32.8 29.9 / 30.2 31.0 / 29.6 29.6 / 29.9 32.0 / 30.2 31.7 38.1 HumanEval 20.7 22.0 22.6 23.2 22.6 20.3 / 19.5 21.6 / 20.3 20.1 / 18.9 24.4 / 20.7 26.8 31.7 Average value 25.1 25.5 25.6 24.9 24.8 21.1 / 21.2 25.0 / 24.3 22.7 / 21.6 27.4 / 23.6 29.5 31.1
4.3 Ablating the Magnitude Principle in LoRA Improvements
We conduct ablation experiments on LLaMA-2-7B touvron2023llama under the NLG setting, focusing on the effect of magnitude gain factor, the choice of basis matrix, and the validation of the magnitude principle.
Magnitude Gain Factor. As shown in Table 4, we first evaluate different values of $Q[r]$ and observe that increasing its value slightly improves performance. We further introduce a “tracking mode” that adjusts $\beta$ in Algorithm 1 based on the initialization magnitudes from reference methods. Specifically, we set $\beta=\sqrt{\frac{\mathbb{V}[B_{\text{ref}}A_{\text{ref}}]}{\mathbb{V}\left[\Phi_{n}\Phi_{m}^{\top}\right]}}$, where $B_{\text{ref}}$ and $A_{\text{ref}}$ are initialization matrix using weight-driven PiSSA ; MiLoRA or data-driven approaches corda ; lora-ga . Under this mode, LoRAM essentially matches the performance of prior methods, confirming that magnitudes govern its performance. We also observe that PiSSA, which selects the top-$r$ singular values, outperforms MiLoRA, which selects the last $r$ singular values. This indicates that leveraging the large dominant singular values enhances performance.
Basis Matrix. We find that the choice of basis matrix generally has limited impact. For instance, replacing the DST basis with a random orthogonal or Gaussian matrix just results in only minor performance degradation. In tracking mode, substituting the SVD-derived basis with DST does not significantly impact performance. A notable exception is LoRA-GA lora-ga , which approximates the full-parameter gradient at initialization. Nonetheless, we emphasize that tracking mode fails not due to incorrectness of magnitude principle, but because the LoRA-GA matrix form maximizes gradient magnitudes of Eq. (2), making it irreplaceable by alternatives and validating magnitude principle [^2] See Proposition 5 in Appendix. We prove LoRA-GA initialization maximizes LoRA gradient magnitude..
Upper Bound of Magnitude Scaling. While increasing update magnitude generally improves performance, the benefit is not unlimited. For example, applying RsLoRA to LoRA-GA yields a clear gain at rank 8, but the improvement diminishes and may even reverse at rank 128. This suggests that magnitude scaling should be applied conservatively at higher ranks, since larger ranks inherently amplify updates, as demonstrated in our Proposition 2. Given that data-driven methods involve costly gradient and SVD computations, we recommend LoRAM with RsLoRA as a more efficient and scalable alternative for accelerating LoRA convergence and performance.
5 Conclusion
In this paper, we explore the magnitude principle of Low-Rank Adaptation (LoRA) and introduce a novel magnitude-driven initialization strategy, LoRAM, that bridges the gap between efficiency and performance. Our work demystifies the prevailing awareness surrounding spectral initialization methods, demonstrating that their success primarily stems from the amplification of weight update magnitudes. By focusing on magnitude regulation as the key driver of convergence, we provide a unified perspective that connects seemingly disparate hyperparameter adjustments, such as learning rate, scaling factor, and initialization schemes, under a single framework.
Limitations and Future Work. Despite the advancements introduced in this work, several challenges remain open for future research. First, LoRAM mimics spectral initialization magnitudes rather than seeking optimal ones; exploring alternative strategies could yield further gains. Additionally, different layers may benefit from tailored magnitude settings, motivating joint optimization with learning rate and rank. Finally, our work does not explicitly address optimization dynamics and convergence properties. These directions remain valuable for advancing parameter-efficient fine-tuning.
Appendix A Related Work
Advances in Low-Rank Adaptation.
LoRA (Hu2021LoRALA, ) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low, ; mao2025survey, ). Existing improvements to LoRA primarily focus on three directions:
Knowledge-Driven Low-Rank Initialization.
Proper initialization critically influences neural network training outcomes glorot2010understanding ; he2015delving . The standard LoRA (Hu2021LoRALA, ) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence. Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA, ; MiLoRA, ; olora, ; SLoRA, ) or task-specific data (lora-one, ; corda, ; lora-ga, ). Most prominent approaches involve Singular Value Decomposition (SVD), which allows flexible control over matrix rank. As a pioneer, PiSSA PiSSA initializes LoRA weights using principal singular components of pre-trained matrices, aligning adaptation directions with the most significant parameter variations to accelerate convergence. Subsequent works corda ; lora-ga ; lora-one ; paischer2024one propose data-driven strategies that incorporate domain knowledge into adapter construction. For instance, LoRA-GA (lora-ga, ) aligns low-rank gradient directions with full fine-tuning counterparts during initialization. While these methods boost performance and retain the core LoRA structure, their theoretical effects on optimization dynamics remain unclear. More importantly, they are less pluggable than LoRA, requiring extra computational pipelines and storage for SVD buffers.
Optimization Dynamics of LoRA. LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial, ). Most existing theoretical studies are limited to simplified and idealized scenarios, such as the lazy-training regime jang2024lora ; malladi2023kernel and infinite-width limitlora+ ; hayou2024impact . A common goal across existing studies involves ensuring feature learning stability to prevent unstable or collapsed training dynamics. For instance,  kalajdzievski2023rank proves that improper scaling factors induce gradient collapse in high-rank adapters, proposing modified scaling mechanisms to stabilize forward and backward propagation. Recent work lora+ ; zhu2024asymmetry ; hayou2024impact reveals critical asymmetries in LoRA: The two low-rank matrices exhibit divergent distinct impacts on optimization trajectories. This asymmetry motivates lora+ to employ distinct learning rates for each matrix. For SVD-based initialization, lora-one investigates principal components derived from single-step full fine-tuning gradients lora-ga , providing convergence guarantees yet under restrictive assumptions. Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA, ).
The intricate nature of LoRA and its numerous advancements motivates us to seek a streamlined principle explaining its empirical success and guiding practical applications. We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank, ; biderman2024lora, ), initialization strategies (PiSSA, ; lora-ga, ), and learning rates (lora+, ), into a cohesive perspective.
Appendix B Lower Bound on Representation Error
Proposition 1 (Lower Bound on Representation Error).
Consider the function class of LoRA-parameterized linear models:
=$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
Let $R(f)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]$ denote the expected regression loss under data distribution $\mathcal{D}$, with squared error loss $\ell(f(x),y)=\|y-f(x)\|^{2}$. Define $f^{*}=\arg\min_{f}R(f)$ as the globally optimal predictor and $f^{*}_{\mathcal{H}}=\arg\min_{f\in\mathcal{H}}R(f)$ as the optimal predictor within $\mathcal{H}$. With loss of generality, we consider the optimal predictor is linear, i.e., $f^{*}(x)=W^{*}x$. Assume that the input covariance matrix is positive definite, i.e., $\Sigma_{\mathcal{D}}=\mathbb{E}_{x\sim\mathcal{D}}[xx^{\top}]\succ 0$, and the magnitudes is limited, satisfying $\alpha r\sqrt{mnM_{1}M_{2}}<\|W^{*}-W_{0}\|_{F}$. Then, the representation error is strictly positive and lower bounded:
$R(f^{*}_{\mathcal{H}})-R(f^{*})\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\left(\|W^{*}-W_{0}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}\right)^{2}>0,$
where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue of the matrix.
Proof.
Due to the linear learnability, we have that $y(x)=f^{*}(x)=W^{*}x$. For any $W$, we have that
$R(W)$ $=\mathbb{E}_{x\sim\mathcal{D}}\left[\|(W-W^{*})x\|^{2}\right]$ (13)
$=\mathbb{E}_{x\sim\mathcal{D}}\left[(W-W^{*})x\left((W-W^{*})x\right)^{\top}\right]$
$=\operatorname{Tr}\left((W-W^{*})\Sigma_{\mathcal{D}}(W-W^{*})^{\top}\right)$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\|W-W^{*}\|_{F}^{2}.$
The last inequality comes from $\lambda_{\min}(\Sigma_{\mathcal{D}})>0$.
Due to $\nu[A]\leq M_{1},\nu[B]\leq M_{2}$, we have that
$\|A\|_{F}^{2}\leq rmM_{1},\ \|B\|_{F}^{2}\leq rnM_{2}.$
Further, for any $A,\ B\in\mathcal{H}$, we have that
$\|W_{0}+\alpha BA-W^{*}\|_{F}$ $\geq\|W_{0}-W^{*}\|_{F}-\|\alpha BA\|_{F}$ (14)
$\geq\|W_{0}-W^{*}\|_{F}-\alpha\|A\|_{F}\|B\|_{F}$
$\geq\|W_{0}-W^{*}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}$
$>0.$
Therefore, we have that
$\quad\ R(f^{*}_{\mathcal{H}})-R(f^{*})$
$=\min_{\nu[A]\leq M_{1},\nu[B]\leq M_{2}}R(W_{0}+\alpha BA)$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\min_{\nu[A]\leq M_{1},\nu[B]\leq M_{2}}\|W_{0}+\alpha BA-W^{*}\|_{F}^{2}$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\left(\|W_{0}-W^{*}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}\right)^{2}>0.$
The first inequality comes from Eq. (13) and the last inequality comes from Eq. (14). ∎
Appendix C Proof of Main Theorems
C.1 Proof of Parameter Scaling Equivalence
Proposition 2 (Parameter Scaling Equivalence).
For LoRA layers defined in Eq. (1), consider decomposing the scaling factor $\alpha=\alpha^{\prime}\alpha_{A}\alpha_{B}$, where $\alpha^{\prime},\alpha_{A},\alpha_{B}\in\mathbb{R}^{+}$. Under the commonly used optimization frameworks with negligible numerical errors, the following parametrization schemes exhibit dynamical equivalence throughout training: For all iterations $t\geq 0$, $\Delta W^{(t)}_{\mathrm{LoRA}}=\Delta\tilde{W}^{(t)}_{\mathrm{LoRA}}$ and $W^{(t)}_{\mathrm{LoRA}}=\tilde{W}^{(t)}_{\mathrm{LoRA}}$, where $\tilde{A}^{(t)}$, $\tilde{B}^{(t)}$ and $\tilde{W}^{(t)}$ represent the re-parameterized versions.
Original SGD Adam
Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
Proof.
Given LoRA’s weight decomposition $W_{\mathrm{LoRA}}=\alpha BA$ where $\alpha>0$, the gradient updates follow:
$\Delta W_{\mathrm{LoRA}}^{(t)}=\alpha(B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}),$ (15)
with learning rates $\eta_{A},\eta_{B}$ for parameters $A$ and $B$, respectively. We define reparameterized parameters $\tilde{A}=\alpha_{A}A$, $\tilde{B}=\alpha_{B}B$, and $\alpha^{\prime}=\alpha/(\alpha_{A}\alpha_{B})$. The LoRA projection becomes:
$W_{\mathrm{LoRA}}=\alpha BA=\alpha^{\prime}\tilde{B}\tilde{A}.$ (16)
This transformation preserves the functional form while redistributing scaling factors.
Proof for SGD
Under initial conditions:
$\tilde{A}^{(0)}$ $=\alpha_{A}A_{\text{init}},$ (17)
$\tilde{B}^{(0)}$ $=\alpha_{B}B_{\text{init}}.$ (18)
Gradients for reparameterized parameters:
$\nabla_{\tilde{A}}L$ $=\alpha^{\prime}\tilde{B}^{\top}\nabla_{W}L,$ (19)
$\nabla_{\tilde{B}}L$ $=\alpha^{\prime}\nabla_{W}L\tilde{A}^{\top}.$ (20)
With learning rates $\eta_{\tilde{A}}=\alpha_{A}^{2}\eta_{A}$, $\eta_{\tilde{B}}=\alpha_{B}^{2}\eta_{B}$ and $t=0$:
$\tilde{A}^{(t+1)}$ $=\tilde{A}^{(t)}-\alpha_{A}^{2}\eta_{A}\nabla_{\tilde{A}}L^{(t)}$ (21)
$=\alpha_{A}A^{(t)}-\alpha_{A}^{2}\eta_{A}(\alpha^{\prime}\alpha_{B}B^{(t)})^{\top}\nabla_{W}L^{(t)}$ (22)
$=\alpha_{A}\left(A^{(t)}-\alpha\eta_{A}(B^{(t)})^{\top}\nabla_{W}L^{(t)}\right)$ (23)
$=\alpha_{A}\left(A^{(t)}-\eta_{A}\nabla_{A}L^{(t)}\right)$ (24)
$=\alpha_{A}A^{(t+1)}.$ (25)
Similarly for $\tilde{B}$:
$\tilde{B}^{(t+1)}$ $=\alpha_{B}\left(B^{(t)}-\eta_{B}\nabla_{B}L^{(t)}\right)=\alpha_{B}B^{(t+1)}.$ (26)
These equations propagate through all $t>0$ via mathematical induction. Compare weight increments:
$\Delta\tilde{W}_{\mathrm{LoRA}}^{(t)}$ $=\alpha^{\prime}[\tilde{B}^{(t+1)}\tilde{A}^{(t+1)}-\tilde{B}^{(t)}\tilde{A}^{(t)}]$ (27)
$=\alpha^{\prime}\alpha_{A}\alpha_{B}[B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}]$ (28)
$=\alpha[B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}]$ (29)
$=\Delta{W}_{\mathrm{LoRA}}^{(t)}.$ (30)
Proof for Adam
Adam maintains exponential moving averages ($m_{t}$, $v_{t}$) for gradients. Under parameter scaling $\tilde{P}=kP$, gradients transform as $\nabla_{\tilde{P}}L=k\nabla_{P}L$. The momentum terms inherit scaling factors:
$m_{\tilde{P}}$ $=\beta_{1}m_{\tilde{P}}+(1-\beta_{1})k\nabla_{P}L,$ (31)
$v_{\tilde{P}}$ $=\beta_{2}v_{\tilde{P}}+(1-\beta_{2})k^{2}(\nabla_{P}L)^{2}.$ (32)
The pratical gradients for Adam:
$\nabla^{\dagger}\tilde{P}$ $=m_{\tilde{P}}/\sqrt{v_{\tilde{P}}+\epsilon}$ (33)
$=[\beta_{1}m_{P}+...]/\sqrt{\beta_{2}v_{P}+...}$ (34)
When setting $\epsilon=0$ for alleviating numerical errors, we have $\nabla^{\dagger}\tilde{P}=\nabla^{\dagger}P$. Setting $\eta_{\tilde{P}}=k\eta_{P}$ cancels scaling factors, preserving update magnitudes. With learning rates $\eta_{\tilde{A}}=\alpha_{A}\eta_{A}$, $\eta_{\tilde{B}}=\alpha_{B}\eta_{B}$ and $t=0$:
$\tilde{A}^{(t+1)}$ $=\tilde{A}^{(t)}-\alpha_{A}\eta_{A}\nabla^{\dagger}_{\tilde{A}}L^{(t)}$ (35)
$=\alpha_{A}A^{(t)}-\alpha_{A}\eta_{A}\nabla^{\dagger}_{A}L^{(t)}$ (36)
$=\alpha_{A}A^{(t+1)}.$ (37)
Similarly for $\tilde{B}$. These equalities propagate through all $t>0$ via mathematical induction.
C.2 Proof of Proposition 2
For the convenience, we split the proof of Proposition 2 into two parts.
Proposition 3 (Parameter Magnitude Dynamics).
Consider LoRA parameters $A^{(t)}$ and $B^{(t)}$ updated with learning rate $\eta$. Assume: $A^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{A}^{2}I)$, $B^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{B}^{2}I)$, $\nabla_{W}L\sim\mathcal{N}(\mathbf{0},\sigma_{L}^{2}I)$, and $\mathbb{E}[\langle A^{(t)},\nabla_{A}L^{(t)}\rangle]=\mathbb{E}[\langle B^{(t)},\nabla_{B}L^{(t)}\rangle]=0$. Under these conditions, the parameter norm vector $\boldsymbol{e}_{t}=\left[\mathbb{E}[\|A^{(t)}\|_{F}^{2}],\mathbb{E}[\|B^{(t)}\|_{F}^{2}]\right]^{T}$ evolve as the following linear dynamical system (left) with closed-form solution (right):
$\boldsymbol{e}_{t+1}=\left(I+\begin{bmatrix}0&\gamma_{A}\\ \gamma_{B}&0\end{bmatrix}\right)\boldsymbol{e}_{t},\quad\boldsymbol{e}_{t}=\begin{bmatrix}\frac{\lambda_{+}^{t}+\lambda_{-}^{t}}{2}&\sqrt{\frac{\gamma_{A}}{\gamma_{B}}}\frac{\lambda_{+}^{t}-\lambda_{-}^{t}}{2}\\ \sqrt{\frac{\gamma_{B}}{\gamma_{A}}}\frac{\lambda_{+}^{t}-\lambda_{-}^{t}}{2}&\frac{\lambda_{+}^{t}+\lambda_{-}^{t}}{2}\end{bmatrix}\boldsymbol{e}_{0}.$ (38)
where $\gamma_{A}=m\eta^{2}\sigma_{L}^{2}$, $\gamma_{B}=n\eta^{2}\sigma_{L}^{2}$ and $\lambda_{\pm}=1\pm\sqrt{\gamma_{A}\gamma_{B}}$. Similarly, the parameter magnitudes $\boldsymbol{\nu}_{t}=\left[\mathbb{E}[\nu[A^{(t)}]],\mathbb{E}[\nu[B^{(t)}]]\right]^{T}$ evolve as a linear dynamical system:
$\boldsymbol{\nu}_{t}=\left(I+\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{t-1}.$
Proof.
Following the gradient descent update rule, the parameter dynamics are:
$\begin{cases}A_{t+1}=A_{t}-\eta\nabla_{A}L_{t}\\ B_{t+1}=B_{t}-\eta\nabla_{B}L_{t}\end{cases},$ (40)
where the gradient relationships $\nabla_{A}L=B^{\top}(\nabla_{W}L)$ and $\nabla_{B}L=(\nabla_{W}L)A^{\top}$ are derived from the LoRA architecture. Expand the Frobenius norm for $A_{t+1}$:
$\|A_{t+1}\|_{F}^{2}=\|A_{t}\|_{F}^{2}-2\eta\langle A_{t},\nabla_{A}L_{t}\rangle+\eta^{2}\|\nabla_{A}L_{t}\|_{F}^{2}.$ (41)
Taking expectations (using independence $\mathbb{E}[\langle A_{t},\nabla_{A}L_{t}\rangle]=0$):
$\mathbb{E}\|A_{t+1}\|_{F}^{2}=\mathbb{E}\|A_{t}\|_{F}^{2}+\eta^{2}\mathbb{E}\|\nabla_{A}L_{t}\|_{F}^{2}.$ (42)
Similarly, for $B_{t+1}$:
$\mathbb{E}\|B_{t+1}\|_{F}^{2}=\mathbb{E}\|B_{t}\|_{F}^{2}+\eta^{2}\mathbb{E}\|\nabla_{B}L_{t}\|_{F}^{2}.$ (43)
Expand the norm of $\nabla_{A}L=B^{\top}(\nabla_{W}L)$:
$\|\nabla_{A}L\|_{F}^{2}=\text{Tr}((\nabla_{A}L)(\nabla_{A}L)^{\top})=\text{Tr}(B^{\top}(\nabla_{W}L)(\nabla_{W}L)^{\top}B).$ (44)
Taking expectations (using gradient entry independence):
$\mathbb{E}\|\nabla_{A}L\|_{F}^{2}=\text{Tr}\left(B^{\top}\mathbb{E}[(\nabla_{W}L)(\nabla_{W}L)^{\top}]B\right)=m\sigma_{L}^{2}\|B\|_{F}^{2}.$ (45)
where $\mathbb{E}[(\nabla_{W}L)(\nabla_{W}L)^{\top}]=m\sigma_{L}^{2}I_{n}$ follows from the i.i.d. assumption on gradient entries.
For $\nabla_{B}L=(\nabla_{W}L)A^{\top}$:
$\mathbb{E}\|\nabla_{B}L\|_{F}^{2}=\text{Tr}\left(A\mathbb{E}[(\nabla_{W}L)^{\top}(\nabla_{W}L)]A^{\top}\right)=n\sigma_{L}^{2}\|A\|_{F}^{2}.$ (46)
Substitute gradient norms results:
$\begin{cases}\mathbb{E}\|A_{t+1}\|_{F}^{2}=\mathbb{E}\|A_{t}\|_{F}^{2}+\eta^{2}m\sigma_{L}^{2}\mathbb{E}\|B_{t}\|_{F}^{2}\\ \mathbb{E}\|B_{t+1}\|_{F}^{2}=\mathbb{E}\|B_{t}\|_{F}^{2}+\eta^{2}n\sigma_{L}^{2}\mathbb{E}\|A_{t}\|_{F}^{2}\end{cases}.$ (47)
Define parameters and state vector:
$\gamma_{A}=\eta^{2}m\sigma_{L}^{2},\quad\gamma_{B}=\eta^{2}n\sigma_{L}^{2},\quad\boldsymbol{e}_{t}=\begin{bmatrix}\mathbb{E}\|A_{t}\|_{F}^{2}\\ \mathbb{E}\|B_{t}\|_{F}^{2}\end{bmatrix}.$ (48)
The recursive system becomes:
$\boldsymbol{e}_{t+1}=\begin{bmatrix}1&\gamma_{A}\\ \gamma_{B}&1\end{bmatrix}\boldsymbol{e}_{t}=\left(I+\begin{bmatrix}0&\gamma_{A}\\ \gamma_{B}&0\end{bmatrix}\right)\boldsymbol{e}_{t}.$ (49)
Similar results for $\boldsymbol{\nu}_{t}$ can be obtained using the same proof techniques.
Proposition 4 (Linearized Dynamics Approximation).
For sufficiently small learning rate $\eta$, the closed-form solution admits the first-order approximation:
$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0},\quad\text{where }\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix}.$ (50)
This yields the update variance expansion under small-$\eta$ regime:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+k_{2}\gamma^{2}t^{2},$ (51)
with $\gamma=\eta^{2}\sigma_{L}^{2}$, $k_{1}=r(m\nu[A^{(0)}]^{2}+n\nu[B^{(0)}]^{2})$, and $k_{2}=rmn\nu[A^{(0)}]\nu[B^{(0)}]$.
Proof.
For a sufficiently small learning rate $\eta$, the product $\gamma_{A}\gamma_{B}=mn\eta^{4}\sigma_{L}^{4}$ is very small. Hence, the eigenvalues
$\lambda_{\pm}=1\pm\sqrt{\gamma_{A}\gamma_{B}}$ (52)
can be approximated via a first-order Taylor expansion. Thus, for small $t$, the closed-form solution for $\boldsymbol{\nu}_{t}$ can be approximated as
$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0}.$ (53)
By definition, we set
$\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\,\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\,\gamma_{A}\sigma_{A}^{2}\end{bmatrix},$ (54)
which agrees with the first-order expansion of the linear system. ∎
Appendix D Analyzing LoRA-GA from Magnitude Principle
Proposition 5 (LoRA-GA initialization maximizes low-rank gradient magnitude).
Given a gradient matrix $\nabla_{W}L\in\mathbb{R}^{n\times m}$ with SVD decomposition $\nabla_{W}L=USV^{\top}$, the optimal rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$ in Eq. (2), under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, are given by:
$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r},$ (55)
where $V_{:,:r}$ and $U_{:,:r}$ contain the first $r$ right and left singular vectors of $\nabla_{W}L$, respectively.
Proof.
We seek to determine the rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$, subject to the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$. From Eq. (2), the LoRA gradient update follows:
$\nabla_{A}L=\nabla_{W}LB^{\top},\quad\nabla_{B}L=\nabla_{W}^{\top}LA^{\top}.$ (56)
By substituting the singular value decomposition of $\nabla_{W}L$, given by
$\nabla_{W}L=USV^{\top},$ (57)
we obtain
$\nabla_{A}L=USV^{\top}B^{\top},\quad\nabla_{B}L=VSU^{\top}A^{\top}.$ (58)
Taking the Frobenius norm and using the cyclic property of the trace, we have
$\|\nabla_{A}L\|_{F}^{2}=\text{tr}(BS^{2}B^{\top}),\quad\|\nabla_{B}L\|_{F}^{2}=\text{tr}(AS^{2}A^{\top}).$ (59)
To maximize these expressions under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, we use the variational characterization of the singular value decomposition. The optimal choice is to align $A$ and $B$ with the dominant singular vectors of $\nabla_{W}L$, leading to
$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r}.$ (60)
This choice ensures that the gradient magnitude is maximized while satisfying the given constraints, completing the proof. ∎
Appendix E Details on Key Formulas
E.1 Weight Update Magnitude
$\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]\approx r\alpha^{2}\eta^{2}\left(\nu[B^{(t)}]\nu[\nabla_{A}L^{(t)}]+\nu[\nabla_{B}L^{(t)}]\nu[A^{(t)}]\right).$
Proof.
Given the LoRA parameter update rule:
$\Delta W^{(t)}_{\mathrm{LoRA}}=\alpha\eta\left(B^{(t)}\nabla_{A}L^{(t)}+\nabla_{B}L^{(t)}A^{(t)}\right)+\mathcal{O}(\eta^{2}),$ (61)
where we retain first-order terms in $\eta$ under small learning rate assumption.
Assumptions:
  • (A1)
    Independence: $B\perp\!\!\!\perp\nabla_{A}L$, $\nabla_{B}L\perp\!\!\!\perp A$
  • (A2)
    Zero-mean initialization: $\mathbb{E}[B_{ij}]=\mathbb{E}[A_{kl}]=0$
  • (A3)
    Spatial homogeneity: $var[B_{ij}]=\nu[B]$, $var[(\nabla_{A}L)_{ij}]=\nu[\nabla_{A}L]$.
For entry $(m,n)$ in $\Delta W^{(t)}_{\mathrm{LoRA}}$:
$var\left[(\Delta W)_{mn}\right]$ $\approx\alpha^{2}\eta^{2}var\left[\sum_{k=1}^{r}\left(B_{mk}(\nabla_{A}L)_{kn}+(\nabla_{B}L)_{mk}A_{kn}\right)\right]$ (62)
$=\alpha^{2}\eta^{2}\left({var\left[\sum_{k}B_{mk}(\nabla_{A}L)_{kn}\right]}+var\left[\sum_{k}(\nabla_{B}L)_{mk}A_{kn}\right]\right).$ (63)
By assumptions (A1)-(A3):
$var\left[\sum_{k}B_{mk}(\nabla_{A}L)_{kn}\right]$ $=\sum_{k}var[B_{mk}]var[(\nabla_{A}L)_{kn}]$ (64)
$=r\nu[B]\nu[\nabla_{A}L].$ (65)
Similarly:
$var\left[\sum_{k}(\nabla_{B}L)_{mk}A_{kn}\right]$ $=r\nu[\nabla_{B}L]\nu[A].$ (66)
Combining these terms:
$\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]$ $\approx\alpha^{2}\eta^{2}\left(r\nu[B]\nu[\nabla_{A}L]+r\nu[\nabla_{B}L]\nu[A]\right)$ (67)
$=r\alpha^{2}\eta^{2}\left(\nu[B]\nu[\nabla_{A}L]+\nu[\nabla_{B}L]\nu[A]\right).$ (68)
where $\nu[B]$, $\nu[A]$ inherit their magnitudes from initialization scheme, and $\nu[\nabla L]$ terms reflect task-specific loss landscape characteristics. ∎
E.2 Spectral Concentration Factor
Let the singular values of the pretrained weight matrix $W$ be $\{s_{i}\}_{i=1}^{\mathcal{R}[W]}$. Define the average of the top-$r$ singular values as
$\mathbb{E}_{r}[s]=\frac{1}{r}\sum_{i=1}^{r}s_{i},$ (69)
and the average of the squares of all singular values as
$\mathbb{E}_{\mathcal{R}[W]}[s^{2}]=\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}.$ (70)
Because the square function is convex, Jensen’s inequality implies
$\mathbb{E}_{r}[s]^{2}\leq\mathbb{E}_{r}[s^{2}],$ (71)
with equality if and only if all $s_{i}$ (for $i=1,\dots,r$) are equal. Moreover, since lower singular values generally contribute less to the overall energy, as $r$ increases the value of $\mathbb{E}_{r}[s]$ decreases relative to $\mathbb{E}_{\mathcal{R}[W]}[s^{2}]$, making $\rho[r]$ a monotonically decreasing function of $r$.
E.3 Variance of $A_{\mathrm{SVD}}$ and $B_{\mathrm{SVD}}$
Recall that the PiSSA initialization is given by
$A_{\mathrm{SVD}}=\sqrt{S_{r}}\,V_{:,:r}^{\top},\quad B_{\mathrm{SVD}}=U_{:,:r}\sqrt{S_{r}},$ (72)
where $S_{r}$ is a diagonal matrix containing the top-$r$ singular values of $W$. Assuming that the columns of $V$ (and similarly, $U$) form an orthonormal basis, the variance of $V_{:,:r}$ (taken elementwise) is approximately $\nu[V_{:r}]\approx\frac{1}{m}$ (or $\frac{1}{n}$ for $U$), since for an orthogonal matrix the energy is uniformly distributed. Thus, the variance of $A_{\mathrm{SVD}}$ can be expressed as:
$\nu(A_{\mathrm{SVD}})=\frac{\mathrm{Tr}(\sqrt{S^{T}_{r}}V_{:,:r}^{T}V_{:,:r}\sqrt{S})}{mr}=\frac{\sum_{i=1}^{r}s_{i}}{mr}=\mathbb{E}_{r}[s]\nu[V_{:r}].$ (73)
To connect this with the variance of $W$, note that
$\nu[W]=\frac{1}{mn}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}.$ (74)
Thus, we can relate $\mathbb{E}_{r}[s]$ to $\nu[W]$ via the spectral concentration factor $\rho[r]$. Incorporating a factor of $n$ to account for the dimensions of $A_{\mathrm{SVD}}$, we obtain:
$\nu(A_{\mathrm{SVD}})=\sqrt{\frac{n\,\rho[r]\,\nu[W]}{m\,\mathcal{R}[W]}}.$ (75)
A similar argument, with the roles of $m$ and $n$ interchanged, leads to the expression for $\nu(B_{\mathrm{SVD}})$.
E.4 Spectral Gain Factor
The dynamics of the LoRA weight update variance are captured by an expression of the form:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+k_{2}\gamma^{2}t^{2},$ (76)
where $k_{1}$ is the linear evolution rate. Substituting the variance expressions derived for $A_{\mathrm{SVD}}$ and $B_{\mathrm{SVD}}$ into the dynamical system (see Eq. (6)), we obtain:
$k_{1}=\frac{\rho[r]\,r\,(m+n)}{\mathcal{R}[W]}\,\nu[W].$ (77)
Defining
$Q[r]\triangleq\frac{\rho[r]\,r}{\mathcal{R}[W]},$ (78)
this expression becomes:
$k_{1}=Q[r](m+n)\nu[W],$ (79)
with the constraint $0\leq Q[r]\leq 1$. When $\nu[W]\sim\mathcal{O}(\min(1/m,1/n))$, the factor $Q[r]$ effectively quantifies the amplification of the weight update magnitude due to the spectral initialization.
Proof.
Recall that
$\rho[r]\triangleq\frac{\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}}{\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}},$ (80)
and the spectral gain factor is defined as
$Q[r]\triangleq\frac{\rho[r]\,r}{\mathcal{R}[W]}.$ (81)
Let $m=\mathcal{R}[W]$. By Jensen’s inequality (or by the Cauchy–Schwarz inequality), we have
$\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}\leq\frac{1}{r}\sum_{i=1}^{r}s_{i}^{2}.$ (82)
Therefore,
$\rho[r]\leq\frac{\frac{1}{r}\sum_{i=1}^{r}s_{i}^{2}}{\frac{1}{m}\sum_{i=1}^{m}s_{i}^{2}}=\frac{m}{r}\cdot\frac{\sum_{i=1}^{r}s_{i}^{2}}{\sum_{i=1}^{m}s_{i}^{2}}\leq\frac{m}{r},$ (83)
since $\frac{\sum_{i=1}^{r}s_{i}^{2}}{\sum_{i=1}^{m}s_{i}^{2}}\leq 1$. Substituting this bound into the definition of $Q[r]$ yields
$Q[r]=\frac{\rho[r]\,r}{m}\leq\frac{m}{r}\cdot\frac{r}{m}=1.$ (84)
This completes the proof that $Q[r]\leq 1$. ∎